OpenStudy (anonymous):

If f(x) = log3 (x + 1), what is f−1(2)? @Salmon

OpenStudy (anonymous):

f-1(x)=3^x -1 f-1(2)=3^2 -1=9-1=8

OpenStudy (alekos):

I get f-1(x) = 3^(x-1) -1 ? salmon can you double check?

OpenStudy (anonymous):

\[y=\log_{3} (x+1)\] converting to exponential form \[3^y=(x+1)\] \[3^y-1=x\] so, \[f^{-1}(x)=x^3-1\]

OpenStudy (isaiah.feynman):

@alekos I think salmon is wrong.

OpenStudy (isaiah.feynman):